If a painter takes 3 hours paint a house, other painter 5 hours, how many hours take both together paint the house?
Answer by ptaylor(2198)

Let x = amount of time it takes both painters, working together, to paint the house
Then, working together, they paint at the rate of 1/x house per hour
First painter paints at the rate of 1/3 house per hour
Other painter paints at the rate of 1/5 house per hour
Together they paint at the rate of 1/3 + 1/5=5/15 + 3/15= 8/15 house per hour
Soooooo
8/15 = 1/x
8x=15
x=1 7/8 hours
CK
(1 7/8=15/8); (15/8)*(8/15)=1 (1 house, that is)
1=1
Hope this helps--ptaylor